OpenStudy (anonymous):

Diff Eq dy/dx=(y^2-y)/cosx, y(o)=1/2

OpenStudy (mathmale):

You might find this problem easier to understand, as I do, if you'd write it in Equation Editor:\[\frac{ dy }{ dx }=\frac{ y^2-y }{ \cos x }\]

OpenStudy (mathmale):

This appears to be a SEPARABLE differential equation. Please separte it, so that all the x-terms are on one side of your equation, and all y-terms are on the other. Then ask yourself: what next?

OpenStudy (mathmale):

@AlexLyu ?? Please become involved in solving this d. e. If you're not sure of what to do next, please ask questions.

OpenStudy (anonymous):

Hi, thanks for the reply, after doing the separable, i got stuck on how to solve the integral 1/cosx.

OpenStudy (mathmale):

Just re-write 1/cos x as sec x. There is an integration formula for sec x.

OpenStudy (accessdenied):

Were you still having trouble with this problem because you bumped it? If so, please tell us your issue! :)